SOLVE :- sin(inverse)2a(sqr)/(1+a(sqr)) + cos(inverse)(1-b(sqr))/(1+b(sqr)) = 2tan(inverse)x

Grade:11

1 Answers

Jayanth Kasaraneni
7 years ago
=sin(inverse)2a(sqr)/(1+a(sqr)) + cos(inverse)(1-b(sqr))/(1+b(sqr)) = 2tan(inverse)x
=2tan(inverse)a+2tan(inverse)b=2tan(inverse)x
=tan(inverse)a+tan(Inverse)b=tan(inverse)x
=tan(inverse)(a+b/1-ab)=tan(inverse)x
=(a+b)/(1-ab)=x